OpenStudy (anonymous):

The Blackbird Oil Company is considering the purchase of 20 new jumbo oil storage tanks. The standard model holds 12,000 gallons. Its dimensions are the size of the similarly shaped jumbo model, that is, the ratio of the dimensions is 4:5. a. How much more storage capacity would the purchase of the twenty jumbo models give Blackbird Oil? b. If jumbo tanks cost 50% more than standard tanks, which tank is a better buy?

OpenStudy (dumbcow):

at a ratio of 4:5, 1 jumbo tank holds 15,000 gallons 4/5 = 12,000/x --> x = 12,000(5/4) = 15,000 20 jumbo tanks --> 15,000*20 = 300,000

OpenStudy (dumbcow):

Let x be cost of 20 standard tanks x + 50%x = x+.5x = 1.5x cost of 20 jumbo tanks is 1.5x 20 standard tanks hold 12,000*20 = 240,000 gal unit price = x/240,000 20 jumbo tanks holds 300,000 unit price = 1.5x/300,000 = x/200,000 x/240,000 < x/200,000 Therefore, the standard tanks are the best value